# Stochastic Population Update Provably Needs An Archive in Evolutionary Multi-objective Optimization

### v1

@@ 4.2 An Archive is Provably Helpful
In the last section, we showed that for NSGA-II solving OneJumpZeroJump, the acceleration brought by SPU is $\Theta((k/e\ln C)^{k-1})$, where $C=e/(p_{s}(1-p_{c}))$, significantly improving the running time $O(\mu n^{k})$ without SPU Doerr and Qu (2023a). Now, we show that using an archive, the result can be further improved.
@@
We prove in Theorem 6 that if using an archive, a population size $\mu\geq 5$ is sufficient to guarantee the same running time bound of NSGA-II with SPU as Theorem 4. Note that Theorem 4 requires $\mu\geq 4(\mu-2k+3)/(1-2p_{s})$, which implies that using an archive can allow a small population size and thus bring speedup. The proof idea is similar to that of Theorem 3, which divides the optimization procedure into three phases, where the first phase aims at finding the two boundary solutions of $F_{I}^{*}$ (i.e., a solution with $k$ 1-bits and a solution with $n-k$ 1-bits), the second phase aims at finding the two extreme Pareto optimal solutions $1^{n}$ and $0^{n}$, and the third phase aims at finding the remaining objective vectors on the Pareto front.
@@
We divide the running process of NSGA-II into three phases. The first phase starts after initialization and finishes until the two boundary solutions of $F_{I}^{*}$ (i.e., a solution with $k$ 1-bits and a solution with $n-k$ 1-bits) are found; the second phase starts after the first phase and finishes when $1^{n}$ and $0^{n}$ are both found; the third phase starts after the second phase and finishes until the whole Pareto front is found.
@@
For the second phase, the analysis is the same as that of the second phase in Theorem 4. Before reaching the extreme Pareto optimal solution $1^{n}$, the boundary Pareto optimal solution with $n-k$ 1-bits is always the Pareto optimal solution with the maximum $f_{1}$ value and thus preserved in the population. Using SPU, the process of reaching $1^{n}$ can be constructed as jumping across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1,n-1]\}$) through $M$ stepping stones, and any failure during the intermediate jumps will result in restarting the process from the solution with $n-k$ 1-bits. By setting proper values of $M$, this process leads to an expected number of generations, $O(n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$.
Finally, we consider the third phase. The analysis of the third phase is just similar to that of Theorem 3, the only difference is that the probability of selecting $1^{n}$ or $0^{n}$ as a parent solution is changed from $1/\mu$ to $1$, because the fair selection selects each solution in the population once (the order of solution is random) to form the parent population. Then, we directly derive that the expected number of generations of phase 3 is $O(n\log n)$.
Combining the three phases, the total expected number of generations is $O(n\log n+n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\}+n\log n)=O(n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$. The equality holds because $k\geq 2$ and $(en\ln C/k)^{k-1}\geq\log n$ as shown in the last paragraph in proof of Theorem 1. For NSGA-II, each generation requires $\mu$ fitness evaluations. Therefore, the expected number of fitness evaluations is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$. Thus, the theorem holds. ∎
@@
Comparing Theorems 4 and 6, we find that if using SPU, the expected running time for NSGA-II solving OneJumpZeroJump with or without an archive is both $O(\mu n^{k}\cdot(e\ln C/k)^{k-1})$, where $C=e/(p_{s}(1-p_{c}))$. The difference is that using an archive allows for a small constant population size $\mu$, which can bring an acceleration factor of $\Theta(n/(1-2p_{s}))$, as $\mu$ is required to be at least $4(n-2k+3)/(1-2p_{s})$ if not using an archive. Compared to SMS-EMOA, using an archive does not achieve exponential acceleration for NSGA-II. The reason is that NSGA-II employs the $(\mu+\mu)$ update mode and fair selection, allowing it to select and explore every solution in each generation, thereby alleviating selection pressure. This results in a smaller constant $C=e/(p_{s}(1-p_{c}))$ independent of $\mu$ in Theorems 4 and 6, compared to $C=e\mu/(p_{s}(1-p_{c}))$ in Theorems 1 and 3. This also implies that NSGA-II achieves a smaller expected running time than SMS-EMOA, suggesting that the $(\mu+\mu)$ mode may be more suitable for SPU than the $(\mu+1)$ mode.
5 Experiments
In this section, we validate our theoretical findings by considering a well-known practical problem, the multi-objective travelling salesman problem (MOTSP) Ribeiro et al. (2002). MOTSP is a classic combinatorial optimization problem, which can be described as: given a network $L=(V,C)$, where $V=\{v_{1},\dots,v_{D}\}$ is a set of $D$ nodes and $C=\{C_{j}:j\in\{1,\dots,m\}\}$ is a set of $m$ cost matrices between nodes $(C_{j}:V\times V)$, the task is to find the Pareto optimal set of Hamiltonian cycles that minimize each of the $m$ cost objectives.
In our experiments, we consider two 100-cities instances (i.e., kroAB and clusAB) from TSPLIB  [^1] https://webia.lip6.fr/~lustt/Research.html Reinelt (1991). The Pareto front size $|\mathcal{PF}|$ of this problem is known, allowing us to set commensurable population sizes $\mu$ for the compared MOEAs. As for evolutionary operators, we use order crossover ($p_{c}=0.9$) and inversion mutation ($p_{m}=0.05$). For NSGA-II, we use fair selection to select each solution in the population once to form the parent population, as we used in the theoretical analysis. As for the survival probability $p_{s}$ for SPU, we set $p_{s}=1/4$ for NSGA-II and $p_{s}=1/2$ for SMS-EMOA. Note that these settings ensure that both algorithms have at least $\mu/2$ solutions directly preserved to the next generation. We consider NSGA-II and SMS-EMOA under four scenarios with varying population sizes: 1) the original algorithms, 2) using only SPU, 3) using only an archive, and 4) using SPU and an archive. Denoting the Pareto front by $\mathcal{PF}$. According to the conditions in Theorems 1 and 4, the population sizes for NSGA-II and SMS-EMOA should be set to $8|\mathcal{PF}|$ and $2|\mathcal{PF}|$, respectively, to preserve the Pareto-optimal solutions found. However, for ease of comparison and to ensure experimental efficiency, we set the population size $\mu=2|\mathcal{PF}|$ for scenarios where MOEAs do not using an archive. For scenarios using an archive, we cannot simply set a constant population size as in solving OneJumpZeroJump to address a practical problem. Therefore, we conservatively set the population size to $|\mathcal{PF}|/4$. For a fair comparison, we use the same $1,000,000$ fitness evaluations for both MOEAs and their variants, and repeat the experiments for $30$ runs. Table 2 and 3 present the results of a widely-used quality indicator, Inverted Generational Distance (IGD) Coello Coello and Reyes Sierra (2004) on the instance clusAB and kroAB. We used IGD since it can measure how well the obtained solution set represents the Pareto front Li et al. (2020).
As can be seen in the tables, using SPU and an archive simultaneously leads to the best results for both SMS-EMOA and NSGA-II, with their IGD value significantly outperforming the other three variants. This confirms our theoretical results that SPU needs an archive to enhance its exploration ability.
Now let us compare between the two algorithms NSGA-II and SMS-EMOA. For the original algorithms, the IGD results of NSGA-II and SMS-EMOA are very similar. When adding SPU to them, it is clear that NSGA-II considerably outperforms SMS-EMOA. This finding confirms our results in Theorems 1 and 4. Furthermore, when considering the case where both SPU and an archive are used, NSGA-II consistently outperforms SMS-EMOA. This aligns with our results in Theorems 3 and 6 that SPU is more suitable for the $(\mu+\mu)$ update mode of NSGA-II compared to the $(\mu+1)$ update mode of SMS-EMOA.
@@ 6 Conclusion
This paper analytically shows that SPU in MOEAs needs an archive to better leverage its exploration ability. We prove that for the well-known NSGA-II and SMS-EMOA solving OneJumpZeroJump, introducing an archive for SPU can address the dilemma of large population size and may provide an extra exponential speedup. The reason is that SPU requires a large population to preserve the best solutions found, while using an archive allows a small population size, which increases the probability of selecting inferior but undeveloped solutions which are more promising in finding better solutions. We also find that the $(\mu+\mu)$ update mode may be more suitable for SPU than the $(\mu+1)$ update mode. Another theoretical contribution of this paper lies in improving the running time bounds for SMS-EMOA and NSGA-II solving OneJumpZeroJump using SPU. These theoretical findings are empirically validated on the practical MOTSP problem. We hope our work may provide some theoretical evidence for the attempts of designing new MOEAs that separate the exploration (via the evolutionary population) and elitist solution preservation (via an external archive), such as in non-elitist or less elitist MOEAs Tanabe and Ishibuchi (2019); Liang et al. (2023b).

### v2

@@ 4.2 An Archive is Provably Helpful
In the last section, we showed that for NSGA-II solving OneJumpZeroJump, the acceleration brought by SPU is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$, significantly improving the running time $O(\mu n^{k})$ without SPU Doerr and Qu (2023a). Now, we show that using an archive, the result can be further improved.
@@
We prove in Theorem 6 that if using an archive, a population size $\mu\geq 5$ is sufficient to guarantee the same running time bound of NSGA-II with SPU as Theorem 4. Note that Theorem 4 requires $\mu\geq 4(\mu-2k+3)/(1-2p_{s})$, which implies that using an archive can allow a small population size and thus bring speedup.
@@
We divide the running process of NSGA-II into three phases. The first phase starts after initialization and finishes until the two boundary solutions of $F_{I}^{*}$ (i.e., a solution with $k$ 1-bits and a solution with nkn-k 1-bits) are found; the second phase starts after the first phase and finishes when $1^{n}$ and $0^{n}$ are both found; the third phase starts after the second phase and finishes until the whole Pareto front is found.
@@
For the second phase, the analysis is the same as that of the second phase in Theorem 4. Before reaching the extreme Pareto optimal solution $1^{n}$, the boundary Pareto optimal solution with nkn-k 1-bits is always the Pareto optimal solution with the maximum $f_{1}$ value and thus preserved in the population. Using SPU, the process of reaching $1^{n}$ can be constructed as jumping across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1,n-1]\}$) through $M$ stepping stones, and any failure during the intermediate jumps will result in restarting the process from the solution with nkn-k 1-bits. By setting proper values of $M$, this process leads to an expected number of generations, $O(n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$.
Finally, we consider the third phase. The analysis of the third phase is just similar to that of Theorem 3. The only difference is that the probability of selecting $1^{n}$ or $0^{n}$ as a parent solution is changed from $1/\mu$ to $1$, because the fair selection selects each solution in the population once (the order of solution is random) to form the parent population. Then, we directly derive that the expected number of generations of phase 3 is $O(n\log n)$.
Combining the three phases, the total expected number of generations is $O(n\log n+n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\}+n\log n)=O(n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e/(p_{s}(1-p_{c}))$. The equality holds because $k\geq 2$ and $(en\ln C/k)^{k-1}\geq\log n$ as shown in the last paragraph in the proof of Theorem 1. For NSGA-II, each generation requires $\mu$ fitness evaluations. Therefore, the expected number of fitness evaluations is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$. Thus, the theorem holds. ∎
Comparing Theorems 4 and 6, we find that if using SPU, the expected running time for NSGA-II solving OneJumpZeroJump with or without an archive is both $O(\mu n^{k}\cdot(e\ln C/k)^{k-1})$, where $C=e/(p_{s}(1-p_{c}))$. The difference is that using an archive allows for a small constant population size $\mu$, which can bring an acceleration factor of $\Theta(n/(1-2p_{s}))$, as $\mu$ is required to be at least $4(n-2k+3)/(1-2p_{s})$ if not using an archive. Compared to SMS-EMOA, using an archive does not lead to an exponential reduction in the upper bound for NSGA-II. The reason is that NSGA-II employs the $(\mu+\mu)$ update mode and fair selection, allowing it to select and explore all solutions in each generation, thereby alleviating selection pressure. This results in a smaller constant $C=e/(p_{s}(1-p_{c}))$ independent of $\mu$ in Theorems 4 and 6, compared to $C=e\mu/(p_{s}(1-p_{c}))$ in Theorems 1 and 3. This also implies that NSGA-II achieves a smaller expected running time than SMS-EMOA, suggesting that the $(\mu+\mu)$ mode may be more suitable for SPU than the $(\mu+1)$ mode.
5 Experiments
In this section, we conduct experiments on OneJumpZeroJump and the well-known multi-objective travelling salesman problem (MOTSP) Ribeiro et al. (2002). Table 4 presents the results of NSGA-II solving the OneJumpZeroJump problem with size $n\in\{10,15,20,25,30\}$ and $k=3$. The settings are: population size $\mu=8$ with an archive, and $\mu=8(n-2k+3)$ without; survival probability $p_{s}=1/2$ when using SPU; and crossover probability $p_{c}=1/2$. Table 5 presents the results of SMS-EMOA solving the OneJumpZeroJump problem with size $n\in\{10,15,20,25,30\}$ and $k=3$. The settings are: population size $\mu=5$ with an archive, and $\mu=2(n-2k+4)$ without; survival probability $p_{s}=1/2$ when using SPU; and crossover probability $p_{c}=1/2$. We present the average number of fitness evaluations over 200 runs for three configurations: SPU only, archive only, and SPU + archive. Results show that using SPU+archive significantly reduces running time compared to using only SPU, and using only archive.
@@ 5 Experiments
size $n$ SPU Archive SPU+Archive
10 5212.48 11733.52 5125.52
15 35653.92 57941.48 25808.24
20 106424.76 183201.24 65398.97
25 239294.88 415550.16 146384.08
30 479311.56 767354.84 284114.12
Table 4: Average number of fitness evaluations over 200 independent runs for NSGA-II solving OneJumpZeroJump with $k=3$.
size $n$ SPU Archive SPU+Archive
10 8522.88 14213.79 5413.985
15 63693.26 66607.01 26275.59
20 214833.25 158766.44 72884.91
25 579966.42 323110.73 140995.59
30 1176418.18 663510.10 262400.4
Table 5: Average number of fitness evaluations over 200 independent runs for SMS-EMOA solving OneJumpZeroJump with $k=3$.
MOTSP is a classic combinatorial optimization problem, which can be described as: given a network $L=(V,C)$, where $V=\{v_{1},\dots,v_{D}\}$ is a set of $D$ nodes and $C=\{C_{j}:j\in\{1,\dots,m\}\}$ is a set of $m$ cost matrices between nodes $(C_{j}:V\times V)$, the task is to find the Pareto optimal set of Hamiltonian cycles that minimize each of the $m$ cost objectives. For the MOTSP, the Pareto front size $|\mathcal{PF}|$ is known, allowing us to set commensurable population sizes $\mu$ for the compared MOEAs. We conduct experiments on two 100-cities instances (i.e., clusAB and kroAB) of MOTSP  [^1] https://webia.lip6.fr/~lustt/Research.html using NSGA-II and SMS-EMOA under four scenarios with varying population sizes: 1) the original algorithms, 2) using only SPU, 3) using only an archive, and 4) using SPU and an archive. For a fair comparison, each scenario used $1,000,000$ fitness evaluations over $30$ runs. Table 2 and 3 presents the results of a widely-used quality indicator, Inverted Generational Distance (IGD) Coello Coello and Reyes Sierra (2004) on the instance clusAB and kroAB, respectively. We used IGD since it can measure how well the obtained solution set represents the Pareto front Li et al. (2020). Table 2 shows that using SPU and an archive simultaneously can lead to the smallest value of IGD for both SMS-EMOA and NSGA-II, and statistically outperforms the other three scenarios. Moreover, with the SPU method, NSGA-II always achieves smaller IGD values than SMS-EMOA. Similar results can be observed with smaller population sizes. Specifically, Table 6 and Table 7) reports the IGD values obtained under reduced population sizes—namely, $\mu=|\mathcal{PF}|/6$, $|\mathcal{PF}|/8$, $|\mathcal{PF}|/10$, and $|\mathcal{PF}|/12$—on the clusAB and kroAB instances, respectively. The results consistently show that the combination of SPU and an archive maintains superior performance.
Only Archive SPU+Archive
$\mu=|\mathcal{PF}|/6$ 1.60e+4 (2.05e+3) 8.29e+3 (4.32e+2)
$\mu=|\mathcal{PF}|/8$ 1.60e+4 (1.43e+3) 8.92e+3 (1.23e+3)
$\mu=|\mathcal{PF}|/10$ 1.69e+4 (2.10e+3) 1.26e+4 (1.08e+3)
$\mu=|\mathcal{PF}|/12$ 1.34e+4 (1.53e+3) 1.14e+4 (1.65e+3)
Table 6: The IGD Coello Coello and Reyes Sierra (2004) results (mean and standard deviation) of the two variants of NSGA-II on the 100-cities MOTSP instance clusAB from TSPLIB Reinelt (1991). For each MOEA, the best mean is highlighted in bold.
Only Archive SPU+Archive
$\mu=|\mathcal{PF}|/6$ 1.83e+4 (2.00e+3) 1.22e+4 (1.77e+3)
$\mu=|\mathcal{PF}|/8$ 1.60e+4 (1.83e+3) 1.16e+4 (1.41e+3)
$\mu=|\mathcal{PF}|/10$ 1.69e+4 (2.10e+3) 1.26e+4 (1.08e+3)
$\mu=|\mathcal{PF}|/12$ 1.58e+4 (1.53e+3) 1.37e+4 (1.83e+3)
Table 7: The IGD Coello Coello and Reyes Sierra (2004) results (mean and standard deviation) of the two variants of NSGA-II on the 100-cities MOTSP instance kroAB from TSPLIB Reinelt (1991). For each MOEA, the best mean is highlighted in bold.
@@ 6 Conclusion
This paper analytically shows that SPU in MOEAs needs an archive to better leverage its exploration ability. We prove that for NSGA-II and SMS-EMOA solving OneJumpZeroJump, introducing an archive for SPU can address the dilemma of large population size and may provide an extra exponential speedup. The reason is that SPU requires a large population to preserve the best solutions found, while an archive enables a small population size, increasing the chance of selecting inferior but undeveloped solutions. We also find that the $(\mu+\mu)$ update mode may be more suitable for SPU than the $(\mu+1)$ update mode. Another contribution lies in improving the running time bounds for SMS-EMOA and NSGA-II solving OneJumpZeroJump using SPU. These theoretical findings are empirically validated on OneJumpZeroJump and the MOTSP. We hope our work may provide some theoretical evidence for the attempts of designing new MOEAs that separate the exploration (via the evolutionary population) and elitist solution preservation (via an external archive), such as in non-elitist or less elitist MOEAs Tanabe and Ishibuchi (2019); Liang et al. (2023b).
Acknowledgements
This work was supported by the National Science and Technology Major Project (2022ZD0116600), the National Science Foundation of China (62276124), and the Fundamental Research Funds for the Central Universities (14380020). Chao Qian is the corresponding author.